# Optimal Multi-Distribution Learning

### v2

@@ 5 Analysis (proof of Theorem 1)
holds for any $1\leq t\leq T$.
@@
Assume lines 1-1 in Algorithm 1 are replaced with some oracle that returns a hypothesis $h^{t}$ satisfying $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$. With probability exceeding $41-\delta/$ 4, the output policy $h^{\mathsf{final}}$ is $\varepsilon$-optimal.
@@
Assume lines 1-1 in Algorithm 1 are replaced with some oracle that returns a hypothesis $h^{t}$ satisfying $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$. With probability at least $21-\delta/$ 2, the sample complexity of Algorithm 1 is bounded by
$O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\left(\log^{5}(k)\log(1/\varepsilon)\log^{2}\left(\frac{k}{\delta\varepsilon}\right)\right)\right).$
5.1 Optimality (proof of Lemma 5)
Let $\mathrm{OPT}:=\min_{h\in\mathcal{H}}\max_{i}L(h,\mathbf{e}_{i})$. Let $v^{t}=L(h^{t},w^{t})-\mathrm{OPT}$ and $f^{t}=\min_{h\in\mathcal{H}}L(h,w^{t})-\mathrm{OPT}\leq 0$. By the assumption that $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$, we have $v^{t}\leq f^{t}+\varepsilon_{1}\leq\varepsilon_{1}$ for any $1\leq t\leq T$.
Let $\delta^{\prime}=\frac{\delta}{T+k+1}$. Note that $\eta\leq 1$ and $|\hat{r}_{i}^{t}|\leq 1$ for any proper $(i,t)$. Direct computation gives that
$\log\left(\frac{\sum_{i=1}^{k}W_{i}^{t+1}}{\sum_{i=1}^{k}W_{i}^{t}}\right)$ $=\log\left(\sum_{i=1}^{k}w_{i}^{t}\exp(\eta\hat{r}^{t}_{i})\right)\leq\log\left(\sum_{i=1}^{k}w_{i}^{t}(1+\eta\hat{r}_{i}^{t}+\eta^{2}(\hat{r}_{i}^{t})^{2})\right)\leq\eta\sum_{i=1}^{k}w_{i}^{t}\hat{r}_{i}^{t}+\eta^{2}.$ (36)
As a result, we have that
$\eta\sum_{t=1}^{T}w^{t}\cdot\hat{r}^{t}$ $\geq\log\left(\sum_{i=1}^{k}W_{i}^{T+1}\right)-T\eta^{2}-\log(k)$
$\geq\max_{i}\log(W_{i}^{T})-T\eta^{2}-\log(k)$
$\geq\eta\max_{i}\sum_{t=1}^{T}\hat{r}_{i}^{t}-T\eta^{2}-\log(k).$ (37)
By dividing both side with $\eta$, we have
$\sum_{t=1}^{T}w^{t}\cdot\hat{r}^{t}\geq\max_{i}\sum_{t=1}^{T}\hat{r}^{t}_{i}-\left(\frac{\log(k)}{\eta}+\eta T\right).$ (38)
By Azuma’s inequality, with probability $1-(k+1)\delta^{\prime}$, we have that
$\left|\sum_{t=1}^{T}w^{t}\cdot\hat{r}^{t}-\sum_{t=1}^{T}L(h^{t},w^{t})\right|\leq 2\sqrt{T\log(1/\delta^{\prime})};$
$\left|\sum_{t=1}^{T}\hat{r}_{i}^{t}-\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})\right|\leq 2\sqrt{T\log(1/\delta^{\prime})}.$
As a result, with probability $1-(k+1)\delta^{\prime}$, it holds that
$\sum_{t=1}^{T}L(h^{t},w^{t})$ $\geq\max_{i}\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})-\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right)$
$\geq-\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right).$ (39)
By the fact that $\varepsilon_{1}=\eta=\frac{1}{100}\varepsilon$ and $T=\frac{20000\log(\frac{k}{\delta^{\prime}\varepsilon})}{\varepsilon^{2}}$, we have that
$\max_{i}\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})$ $\leq T\mathrm{OPT}+\sum_{t=1}^{T}v^{t}+\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right)$
$\leq T\mathrm{OPT}+T\varepsilon_{1}+\left(\frac{\log(k)}{\eta}+\eta T+2\sqrt{T\log(1/\delta^{\prime})}\right)$
$\leq T\mathrm{OPT}+T\varepsilon.$ (40)
As a result, we obtain that
$\max_{i}L(h^{\mathsf{final}},\mathbf{e}_{i})=\max_{i}\frac{1}{T}\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})\leq\mathrm{OPT}+\varepsilon.$ (41)
The proof is finished by noting that $\delta^{\prime}=\frac{\delta}{(T+k+1)}$.
5.2 Sample complexity (proof of Lemma 6)
Recall that $\{w^{t}\}_{t=1}^{T+1}$ is the trajectory by running Algorithm 1. Also recall that $\overline{w}^{T}_{i}=\max_{1\leq t\leq T}w_{i}^{t}$ for all $i\in[k]$. We start with the lemma below to bound the sample complexity by $\|\overline{w}^{T}\|_{1}$.
@@ 5.2 Sample complexity (proof of Lemma 6)
The sample complexity of Algorithm 1 is bounded by $O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\|\overline{w}^{T}\|_{1}\right)$.
@@
Recall the definition of $\check{w}^{t}$ and $\overline{w}^{t}$ in Algorithm 1. It is easy to observe that $\check{w}^{t}\leq\overline{w}^{t}.$ Then the total number of samples is bounded by
$O\left(T_{1}\cdot\|\overline{w}^{T}\|_{1}+T(k\|\overline{w}^{T}\|_{1}+k)\right)=O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\|\overline{w}^{T}\|_{1}\right).$
On the other hand, we have the lemma below to bound $\|\overline{w}^{T}\|_{1}$.
@@
Assume Line 1-1 in Algorithm 1 is replaced by some oracle which returns a hypothesis $h^{t}$ satisfies that $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$. With probability $41-\delta/$ 4, the sample complexity of Algorithm 1 is bounded by
$O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\left(\log^{5}(k)\log(1/\varepsilon)\log^{2}\left(\frac{k}{\delta\varepsilon}\right)\right)\right).$
The proof of Lemma 8 is rather involved. We postpone the proof to Appendix C.1 due to space limitation.
By Lemma 7 and Lemma 8, with probability $21-\delta/$ 2, the sample complexity of Algorithm 1 is bounded by
$O\left(\frac{d\log\left(\frac{d}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\left(\log^{5}(k)\log(1/\varepsilon)\log^{2}\left(\frac{k}{\delta\varepsilon}\right)\right)\right).$
The proof is completed.
6 Extension: learning Rademacher classes
In this section, we study how to adapt our algorithm and theory to accommodate MDL for Rademacher classes.
1 input: $k$ data distributions $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{k}\}$, hypothesis class $\mathcal{H}$, target accuracy level $\varepsilon$, target success rate $1-\delta$, the constant $\{C_{n}\}_{n\geq 1}$ in Assumption 10.
2 hyper-parameters: stepsize $\eta=\frac{1}{100}\varepsilon$, number of rounds $T=\frac{20000\log\left(\frac{k}{\delta}\right)}{\varepsilon^{2}}$, auxiliary accuracy level $\varepsilon_{1}=\frac{1}{100}\varepsilon$, auxiliary sub-sample-size $T_{1}=\min\left\{t\geq\frac{400\left(k\log(\frac{k}{\varepsilon_{1}})+d\log\big{(}(\frac{kd}{\varepsilon_{1}})+\log(\frac{1}{\delta})\big{)}\right)}{\varepsilon_{1}^{2}}\Big{|}C_{t}\leq\frac{\epsilon_{1}}{1200}\right\}$.
@@ 6 Extension: learning Rademacher classes
6       set $w^{t}=[w^{t}_{i}]_{1\leq i\leq k}$ and $\widehat{w}^{t}=[\widehat{w}^{t}_{i}]_{1\leq i\leq k}$, where $w^{t}_{i}\leftarrow\frac{W^{t}_{i}}{\sum_{j}W^{t}_{j}}$ and $\widehat{w}^{t}_{i}\leftarrow\widehat{w}^{t-1}_{i}$ for all $i\in[k]$.
@@ 10                   $n_{i}^{t}\leftarrow\left\lceil T_{1}\widehat{w}^{t}_{i}\right\rceil$;
$\widehat{L}^{t}(h,w^{t}):=\sum_{i=1}^{k}\frac{w^{t}_{i}}{\check{n}_{i}^{t}}\cdot\sum_{j=1}^{\check{n}_{i}^{t}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}$ (42)
with $\check{n}_{i}^{t}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$, and $(x_{i,j},y_{i,j})$ being the $j$-th datapoint from $\mathcal{D}_{i}$ in $\mathcal{S}$.
@@
             draw $\lceil k\overline{w}^{t}_{i}\rceil$ independent samples — denoted by $\big{\{}(x^{t}_{i,j},y^{t}_{i,j})\big{\}}_{j=1}^{\lceil k\overline{w}_{i}^{t}\rceil}$ — from $\mathcal{D}_{i}$, and set
@@
Algorithm 3 Hedge for Multi-distribution Learning on Rademacher Class ($\mathtt{MDL-Hedge-Rademacher}$)
6.1 Preliminaries: Rademacher complexity
Let us first introduce the formal definition of the Rademacher complexity; more detailed introduction can be found in Shalev-Shwartz and Ben-David, (2014).
Definition 9 (Rademacher complexity).
Given a distribution $\mathcal{D}$ supported on $\mathcal{Z}\coloneqq\mathcal{X}\times\mathcal{Y}$ and a positive integer $n$, the (average) Rademacher complexity is defined as
Next, we would like to make an assumption on the Rademacher complexity of mixtures of distributions. Denoting by $\mathcal{D}(w)$ the mixed distribution
$\mathcal{D}(w)\coloneqq\sum_{i=1}^{k}w_{i}\mathcal{D}_{i}$ (44)
for any probability vector $w\in\Delta(k)$, we can state our assumption as follows.
Assumption 10.
For each $n\geq 1$, there exists a universal constant $C_{n}>0$ such that
$\mathsf{Rad}_{n}\big{(}\mathcal{D}(w)\big{)}\leq C_{n}$ (45)
holds for all $w\in\Delta^{k}$.
Remark 1.
One might raise a question about Assumption 10: Can we only assume $\mathrm{Rad}_{n}(\mathcal{D}_{i})\leq C_{n}$ for $i\in[k]$ without incuring a worse sample complexity? The answer is negative. Actually, the Rademacher complexity $\mathsf{Rad}_{n}(\mathcal{D}(w))$ is not convex in $w$, where we fail to use $\max_{i}\mathsf{Rad}_{n}(\mathcal{D}_{i})$ to bound $\max_{w\in\Delta(k)}\mathsf{Rad}_{n}(\mathcal{D}(w))$. We refer to Appendix D.4 for more details.
It is well known that $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$ implies Assumption 10 holds with $C_{n}=\sqrt{\frac{2d\log(en/d)}{n}}$ (Mohri et al.,, 2018).
To facilitate our analysis in this section, we find it helpful to introduce the notion of the weighted Rademacher complexity as follows.
Definition 11 (Weighted Rademacher complexity).
$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}(\mathcal{D})\coloneqq\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\right]\right],$ (46)
The weighted Rademacher complexity defined above satisfies an important property below.
@@ 6.1 Preliminaries: Rademacher complexity
In addition, the following allows us to bound the weighted Rademacher complexity under Assumption 10.
Lemma 13.
Consider any $\{n_{i}\}_{i=1}^{k}$ obeying $n_{i}\geq 12\log(2k)$ for each $i\in[k]$. By taking $w\in\Delta^{k}$ with $w_{i}=\frac{n_{i}}{\sum_{l=1}^{k}n_{l}}$, one has
6.2 Algorithm and sample complexity
We present in Algorithm 3 to learn a Rademacher class under multi-distribution tasks. Algorithm 3 is also a Hedge-like algorithm to learn a convex (concave) game. The major difference from Algorithm 1 lies in the subroutine to learn $h^{t}$ (See lines 3-3 in Algorithm 3). More precisely, to compute the estimator $\widehat{L}^{t}(h,w^{t})$ for $L(h,w^{t})$, instead of using the first $n_{i}^{t}$ samples from $\mathcal{D}_{i}$ for each $i\in[k]$, we choose to use the first $\check{n}_{i}^{t}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$ samples from $\mathcal{D}_{i}$ for each $i$.
Formally, we have the result below.
Theorem 14.
Suppose Assumption 10 holds. By running Algorithm 3, with probability $1-\delta$, the output $h^{\mathsf{final}}$ satisfies that
$\max_{i}L(h^{\mathsf{final}},\mathbf{e}_{i})=\max_{i}\frac{1}{T}\sum_{t=1}^{T}L(h^{t},\mathbf{e}_{i})\leq\min_{h\in\mathcal{H}}\max_{i}L(h,\mathbf{e}_{i})+\varepsilon.$
Meanwhile, the sample complexity of Algorithm 3 is bounded by
$O\left(T_{\epsilon}\cdot\log^{5}(k)\log\left(\frac{1}{\varepsilon}\right)\log^{2}\left(\frac{k}{\delta\varepsilon}\right)\right)$
where $T_{\epsilon}$ is defined as
$T_{\epsilon}=\min\left\{t\geq\frac{400\left(k\log(\frac{k}{\varepsilon_{1}})+d\log\big{(}(\frac{kd}{\varepsilon_{1}})+\log(\frac{1}{\delta})\big{)}\right)}{\varepsilon_{1}^{2}}\Big{|}C_{t}\leq\frac{\epsilon_{1}}{1200}\right\}.$
In the case $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$, we have $C_{n}\leq\sqrt{\frac{2d\log(en/d)}{n}}$, which implies that $T_{\epsilon}=\tilde{O}\left(\frac{d+k}{\epsilon^{2}}\right)$ and a sample complexity bound of $\tilde{O}\left(\frac{d+k}{\epsilon^{2}}\right)$.
Proof of Theorem 14.
By Lemma 5 and Lemma 6, it suffices to show that by running Algorithm 3 with probability $41-\delta/$ 4, $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ for any $1\leq t\leq T$. Formally, we have the lemma below.
Lemma 15.
Suppose Assumption 10 holds. By running Algorithm 3, with probability $41-\delta/$ 4, it holds that

### v3

@@ 5 Analysis for VC classes (proof of Theorem 1)
holds for all $1\leq t\leq T$, where $h^{t}$ (resp. $w^{t}$) is the hypothesis (resp. weight vector) computed in round $t$ of Algorithm 1.
Proof.
See Appendix B.1. ∎
Next, assuming that (29) holds, we can resort to standard analysis for the Hedge algorithm to demonstrate the quality of the final output $h^{\mathsf{final}}$.
@@
Suppose that lines 1-1 in Algorithm 1 are replaced with some oracle that returns a hypothesis $h^{t}$ satisfying $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$. With probability exceeding $41-\delta/$ 4, the hypothesis $h^{\mathsf{final}}$ output by Algorithm 1 is $\varepsilon$-optimal in the sense that
$\max_{1\leq i\leq k}L(h^{\mathsf{final}},\mathrm{e}^{\mathsf{basis}}_{i})\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}L(h,\mathrm{e}^{\mathsf{basis}}_{i})+\varepsilon.$ (30)
Here, we recall that $\mathrm{e}^{\mathsf{basis}}_{i}$ indicates the $i$-th standard basis vector.
Proof.
See Appendix B.2. ∎
Taking Lemma 1 and Lemma 2 together, one can readily see that Algorithm 1 returns an $\varepsilon$-optimal randomized hypothesis $h^{\mathsf{final}}$ with probability at least $21-\delta/$ 2. The next step then lies in bounding the total number of samples that has been collected in Algorithm 1. Towards this end, recall that $\overline{w}^{T}_{i}=\max_{1\leq t\leq T}w_{i}^{t}$ for each $i\in[k]$. Recognizing that $\widehat{w}^{t}_{i}\leq\overline{w}^{t}_{i}$ for each $t\in[T]$ and $i\in[k]$, we can bound the total sample size by
(sample $(\text{sample size})\qquad T_{1}\sum_{i=1}^{k}\widehat{w}_{i}^{T}+k+T\bigg{(}k\sum_{i=1}^{k}\overline{w}_{i}^{T}+k\bigg{)}$ $\leq\left(T_{1}\|\overline{w}^{T}\|_{1}+kT\|\overline{w}^{T}\|_{1}\right)+k(T+1)$
$\lesssim\frac{d\log\left(\frac{kd}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\|\overline{w}^{T}\|_{1},$ (31)
where the last relation follows from our choices $T\asymp\frac{\log(k/\delta)}{\varepsilon^{2}}$ and $T_{1}\asymp\frac{k\log(k/\varepsilon)+d\log(kd/\varepsilon)+\log(1/\delta)}{\varepsilon^{2}}$ (cf. line 1 of Algorithm 1) and the basic property that $\|\overline{w}^{T}\|_{1}\geq\sum_{i}w_{i}^{1}=1$. Consequently, everything then comes down to bounding $\|\overline{w}^{T}\|_{1}$, for which we resort to the following lemma.
@@
Assume that lines 1-1 in Algorithm 1 are replaced with some oracle which returns a hypothesis $h^{t}$ satisfies that $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ in the $t$-th round for each $1\leq t\leq T$. With probability at least $41-\delta/$ 4, the quantity $\|\overline{w}^{T}\|_{1}$ is bounded above by
$\|\overline{w}^{T}\|_{1}\leq O\left(\log^{8}\left(\frac{k}{\delta\varepsilon}\right)\right).$
It is noteworthy that the proof of Lemma 3 is the most technically challenging part of the analysis; we postpone this proof to Appendix B.3.
Combining Lemma 3 with (31) immediately reveals that, with probability at least $1-\delta$, the sample complexity of Algorithm 1 is bounded by
$O\left(\frac{d\log\left(\frac{kd}{\varepsilon}\right)+k\log\left(\frac{k}{\delta\varepsilon}\right)}{\varepsilon^{2}}\cdot\log^{8}\left(\frac{k}{\delta\varepsilon}\right)\right),$
as claimed in Theorem 1. It remains to prove the above key lemmas, which we postpone to Appendix B.
6 Necessity of randomization
Given that the best-known sample complexities prior to our work were derived for algorithms that either output randomized hypotheses or invoke majority votes, awasthi2023sample raised the question about how the sample complexity is impacted if only deterministic (or “proper”) hypotheses are permitted as the output of the learning algorithms. As it turns out, the restriction to deterministic hypotheses substantially worsens the sample efficiency, as revealed by the following theorem.
Theorem 2.
  • a hypothesis class $\mathcal{H}$ containing at most $kN_{0}+1$ hypothesis,
such that: for any algorithm $\mathcal{G}$ that finds $h\in\mathcal{H}$ obeying
with probability exceeding $3/4$, its sample size — denoted by $M(\mathcal{G})$ — must exceed $\mathbb{E}[M(\mathcal{G})]\geq\frac{dk}{240000\varepsilon^{2}}$.
In words, the sample complexity when a deterministic output is required scales at least with $\Omega(\frac{dk}{\varepsilon^{2}}\big{)}$, which is considerably larger than the sample complexity $\widetilde{O}\big{(}\frac{d+k}{\varepsilon^{2}}\big{)}$ achievable via Algorithm 1 (which outputs a randomized hypothesis). The proof of this theorem is deferred to Appendix D.
7 Extension: learning Rademacher classes
In this section, we adapt our algorithm and theory to accommodate MDL for Rademacher classes. Note that when learning Rademacher classes, the hypotheses in $\mathcal{H}$ do not need to be binary-valued.
7.1 Preliminaries: Rademacher complexity
Let us first introduce the formal definition of the Rademacher complexity; more detailed introduction can be found in, e.g., shalev2014understanding.
Definition 2 (Rademacher complexity).
Given a distribution $\mathcal{D}$ supported on $\mathcal{Z}\coloneqq\mathcal{X}\times\mathcal{Y}$ and a positive integer $n$, the Rademacher complexity is defined as
Next, we would like to make an assumption concerning the Rademacher complexity of mixtures of distributions. Denoting by $\mathcal{D}(w)$ the mixed distribution
$\mathcal{D}(w)\coloneqq\sum_{i=1}^{k}w_{i}\mathcal{D}_{i}$ (34)
for any probability vector $w=[w_{i}]_{1\leq i\leq k}\in\Delta(k)$, we can state our assumption as follows.
Assumption 1.
For each $n\geq 1$, there exists a quantity $C_{n}>0$ (known to the learner) such that
$C_{n}\geq\sup_{w\in\Delta(k)}\mathsf{Rad}_{n}\big{(}\mathcal{D}(w)\big{)}.$ (35)
For instance, if the hypothesis class $\mathcal{H}$ obeys $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$, then it is well-known that Assumption 1 holds with the choice (see, e.g., mohri2018foundations)
$C_{n}=\sqrt{\frac{2d\log(en/d)}{n}}.$
Remark 3.
One might raise a natural question about Assumption 1: can we use $\widetilde{C}_{n}\coloneqq\max_{i\in[k]}\mathsf{Rad}_{n}(\mathcal{D}_{i})$ instead of ${C}_{n}$ without incuring a worse sample complexity? The answer is, however, negative. In fact, the Rademacher complexity $\mathsf{Rad}_{n}(\mathcal{D}(w))$ is not convex in $w$, and hence we fail to use $\max_{i}\mathsf{Rad}_{n}(\mathcal{D}_{i})$ to bound $\max_{w\in\Delta(k)}\mathsf{Rad}_{n}(\mathcal{D}(w))$. The interested reader is referred to Appendix E.4 for more details.
To facilitate analysis, we find it helpful to introduce another notion called weighted Rademacher complexity.
Definition 3 (Weighted Rademacher complexity).
$\widetilde{\mathsf{Rad}}_{\{n_{i}\}_{i=1}^{k}}\coloneqq\mathop{\mathbb{E}}\limits_{\{z_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\mathop{\mathbb{E}}\limits_{\{\sigma_{i}^{j}\}_{j=1}^{n_{i}},\forall i\in[k]}\left[\frac{1}{\sum_{i=1}^{k}n_{i}}\max_{h\in\mathcal{H}}\sum_{i=1}^{k}\sum_{j=1}^{n_{i}}\sigma_{i}^{j}\ell\big{(}h,z_{i}^{j}\big{)}\right]\right],$ (36)
The following two lemmas provide useful properties about the weighted Rademacher complexity.
@@ 7.1 Preliminaries: Rademacher complexity
See Appendix E.2.∎
@@
Consider any $\{n_{i}\}_{i=1}^{k}$ obeying $n_{i}\geq 12\log(2k)$ for each $i\in[k]$. By taking $w\in\Delta(k)$ with $w_{i}=\frac{n_{i}}{\sum_{l=1}^{k}n_{l}}$, one has
Proof.
See Appendix E.3.∎
7.2 Algorithm and sample complexity
We are now positioned to introduce our algorithm that learns a Rademacher class in the presence of multiple distributions, which is also based on a Hedge-type strategy to learn a convex-concave game; see Algorithm 3 for full details. Its main distinction from Algorithm 1 lies in the subroutine to learn $h^{t}$ (see lines 3-3 in Algorithm 3) as well as the choice of $T_{1}$ (see line 3 in Algorithm 3). More precisely, to compute the estimator $\widehat{L}^{t}(h,w^{t})$ for $L(h,w^{t})$, instead of using the first $n_{i}^{t}$ samples from $\mathcal{D}_{i}$ for each $i\in[k]$, we choose to use the first
$n_{i}^{t,\mathsf{rad}}=\min\big{\{}\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\big{\}}$
samples from $\mathcal{D}_{i}$ for each $i$, where $T_{1}$ is taken to be
$T_{1}=\min\Big{\{}t\geq\frac{4000(k\log(k/\varepsilon_{1})+\log(1/\delta))}{\varepsilon_{1}^{2}}\,\Big{|}\,C_{t}\leq\frac{\varepsilon_{1}}{4800}\Big{\}}.$ (38)
Here, $T_{1}$ needs to take advantage of the quantities $\{C_{n}\}$ (cf. Assumption 1) that upper bound the associated Rademacher complexity.
1 input: $k$ data distributions $\{\mathcal{D}_{1},\mathcal{D}_{2},\ldots,\mathcal{D}_{k}\}$, hypothesis class $\mathcal{H}$, target accuracy level $\varepsilon$, target success rate $1-\delta$, quantities $\{C_{n}\}_{n\geq 1}$ as in Assumption 1.
2 hyper-parameters: stepsize $\eta=\frac{1}{100}\varepsilon$, number of rounds $T=\frac{20000\log\left(\frac{k}{\delta}\right)}{\varepsilon^{2}}$, auxiliary accuracy level $\varepsilon_{1}=\frac{1}{100}\varepsilon$, auxiliary sub-sample-size $T_{1}=\min\Big{\{}t\geq\frac{4000(k\log(k/\varepsilon_{1})+\log(1/\delta))}{\varepsilon_{1}^{2}}\,\Big{|}\,C_{t}\leq\frac{\varepsilon_{1}}{4800}\Big{\}}$.
@@ 10                   $n_{i}^{t}\leftarrow\left\lceil T_{1}\widehat{w}^{t}_{i}\right\rceil$;
$\widehat{L}^{t}(h,w^{t}):=\sum_{i=1}^{k}\frac{w^{t}_{i}}{n_{i}^{t,\mathsf{rad}}}\cdot\sum_{j=1}^{n_{i}^{t,\mathsf{rad}}}\ell\big{(}h,(x_{i,j},y_{i,j})\big{)}$ (39)
with $n_{i}^{t,\mathsf{rad}}=\min\{\left\lceil T_{1}w_{i}^{t}+12\log(2k)\right\rceil,T_{1}\}$ and $(x_{i,j},y_{i,j})$ being the $j$-th datapoint from $\mathcal{D}_{i}$ in $\mathcal{S}$.
@@
16             draw $\lceil k\overline{w}^{t}_{i}\rceil$ independent samples — denoted by $\big{\{}(x^{t}_{i,j},y^{t}_{i,j})\big{\}}_{j=1}^{\lceil k\overline{w}_{i}^{t}\rceil}$ — from $\mathcal{D}_{i}$, and set
@@
Algorithm 3 Hedge for multi-distribution learning on Rademacher Classes ($\mathtt{MDL}\text{-}\mathtt{Hedge}\text{-}\mathtt{Rad}$)
Equipped with this algorithm, we are ready to establish the following theoretical guarantees.
Theorem 3.
Suppose Assumption 1 holds. With probability at least $1-\delta$, the output $h^{\mathsf{final}}$ returned by Algorithm 3 satisfies
$\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i},h^{\mathsf{final}}}\left[\ell\big{(}h^{\mathsf{final}},(x,y)\big{)}\right]\leq\min_{h\in\mathcal{H}}\max_{1\leq i\leq k}\mathop{\mathbb{E}}\limits_{(x,y)\sim\mathcal{D}_{i}}\big{[}\ell\big{(}h,(x,y)\big{)}\big{]}+\varepsilon.$ (40)
In particular, the total number of samples collected by Algorithm 3 is bounded by
$\left(\frac{k}{\varepsilon^{2}}+\min\big{\{}n\mid C_{n}\leq c_{1}\varepsilon\big{\}}\right)\,\mathrm{poly}\log\Big{(}k,\frac{1}{\varepsilon},\frac{1}{\delta}\Big{)}$
with probability exceeding $1-\delta$, where $c_{1}>0$ is some sufficiently small constant, and $T_{1}$ is defined in (38).
In contrast to the VC classes, the sample complexity for learning Rademacher classes entails a term related to the Rademacher complexity — namely, $\min\big{\{}n\mid C_{n}\leq c_{1}\varepsilon\big{\}}$ — as opposed to the term $d/\varepsilon^{2}$ concerning the VC-dimension. When it comes to the special case where $\mathsf{VC}\text{-}\mathsf{dim}(\mathcal{H})\leq d$, taking $C_{n}\leq\sqrt{\frac{2d\log(en/d)}{n}}$ in (38) leads to $T_{1}=\widetilde{O}\left(\frac{d+k}{\varepsilon^{2}}\right)$, which recovers the sample complexity bound of $\widetilde{O}\left(\frac{d+k}{\varepsilon^{2}}\right)$ derived in Theorem 1 for VC classes.
Proof of Theorem 3.
In view of Lemma 2 and Lemma 3, it boils down to showing that running Algorithm 3 results in $L(h^{t},w^{t})\leq\min_{h\in\mathcal{H}}L(h,w^{t})+\varepsilon_{1}$ for any $1\leq t\leq T$, a property that holds with probability at least $41-\delta/$ 4. To accomplish this, we have the lemma below.
@@
Suppose Assumption 1 holds. With probability at least $41-\delta/$ 4, the iterates of Algorithm 3 satisfy